There are two countries in information in the table about growth and productivity to answer the questions. Round all numerical answers to two decimal places. he world: Funland and Sunhausland. Both countries currently have a GDP per capita of 1. Use the Country Funland Growth rate of GDP per capita 0.100 0.012 Sunhausland What is GDP per cupita in Funland in 17 years? Funlands GDP: $ What is GDP per capita in Sunhausland in 17 years? Sunhauslands GDP: $ In 17 years, Funland grows how many times more than Sunhausland? times Sunhauslands growth Funlands growth:
Answer #1

1) GDP per capita in Funland in 17 years = 1 x (1 + 0.1)17 = 5.05

2) GDP per capita in Sunhausland in 17 years = 1 x (1 + 0.012)17 = 1.22

3) Funland's growth/Sunhausland's growth = 5.05/1.22 = 4.13